# On Modeling and Solving the Boltzmann Equation

### v1

@@ 4.1 Numerical Results and Contributions
The ADO formulation for problems in one-dimensional geometry was developed in this section in the context of the RTE. In addition to the works already cited, the methodology has been applied to a wide range of problems in various areas of transport theory by different authors. It is worth listing here the treatment of models which include polarization effects (Barichello and Siewert, 1999a) of interest in astrophysics, and that, according to our knowledge, has not been solved before; the challenging searchlight problem (Barichello and Siewert, 2000a); ADO for different geometries (Barichello et al., 2002b); medical applications (Groff et al., 2021); the key issues of criticality (Da Cunha et al., 2022) and safeguards (Barichello and Pázsit, 2025) in nuclear engineering; to mention just a few. Later in this work, applications in the field of rarefied gas dynamics will be discussed. We do not list numerical results generated by the solution here. They can be checked in the listed references. However, a central point to note is that in all these applications, the ADO solution proved to be concise and accurate, requiring a lower computational cost compared to what was available in the literature. The natural challenge, considering more realistic problems, was to extend the idea to multidimensional problems, which are the subject of the next section.
@@ 5 The ADO Solution for Two-dimensional Geometry Problems
Purely numerical methods dominate simulations for multidimensional transport models. The use of analytical techniques is limited and very challenging. However, they are of great importance for generating benchmark solutions. One of the first issues regarding discrete-ordinate methods, which already differ greatly from one-dimensional problems, is the definition of numerical quadrature schemes for integration on the unit sphere. Madsen Madsen (1971) identified the truncation error in the approximation of the integral term of the equation as a fundamental point in the convergence of the discrete-ordinate solution.
@@ 5.1 Numerical Quadrature Schemes
The choice of an appropriate quadrature scheme to approximate the integral term in Eq. (1) is a relevant issue, in particular for multidimensional problems. Traditionally the quadratures were required to preserve the reflective and rotational symmetry properties and, in this context, the classical Level Symmetric ($LQ_{N}$) quadrature scheme (Lewis and Miller, 1984) is widely used for dealing with multidimensional discrete ordinates problems (see Fig (3)). It is well known, however, that such scheme is limited to order $N\leq 20$. The number of directions per octant is given by $M=N(N+2)/8$.
In our studies, we have additionally considered Legendre-Chebyshev triangular ($P_{N}T_{N}S_{N}$) and quadrangular ($P_{N}T_{N}$) schemes (Longoni and Haghighat, 2001; Cacuci, 2010) and Quadruple Range ($QR$) schemes (Abu-Shumays, 1977), which enable generate numerical results for much higher quadrature orders, even if some (rotational) symmetry properties are relaxed.
The Legendre-Chebyshev quadrature schemes utilize a one-dimensional Gaussian quadrature to treat the polar variable $\theta$ and an equally weighted Chebyshev quadrature for the azimuthal variable $\phi$ (Longoni and Haghighat, 2001; Cacuci, 2010). Thus, in a three-dimensional geometry, the ordinates $\xi=\cos\theta$ are given by the roots of the Legendre polynomials ($P_{N}$). The azimuthal angular variable in each level is defined by the roots of the Chebyshev polynomials ($T_{N}$) of the first kind. The Legendre-Chebyshev quadrangular scheme ($P_{N}T_{N}$) (Cacuci, 2010) defines $M=(N^{2}/4)$ discrete directions per octant, where for each level $\xi_{i}$ the corresponding directions have the same weight (Fig. (4). For the two-dimensional case, one coordinate is disregarded. Still, the corresponding quadrature weights are given in terms of $w_{i}$, the weights of the Gauss-Legendre quadrature.
Figure 4: Legendre-Chebishev Quadrangular Scheme
For the Legendre-Chebyshev Triangular scheme ( $P_{N}T_{N}S_{N}$) (Longoni and Haghighat, 2001), on the other hand, the $M=N(N+2)/8$ discrete directions per octant follow a pattern analogous to the Level Symmetric quadrature scheme. The set of azimuthal angles for each level $\xi_{i}$ is defined as
@@
The Quadruple Range quadrature ($QR$) scheme, proposed by Abu-Shumays (1977) , is a product quadrature. As a general description, the polar angle $\theta$ is approximated by Gauss-Christoffel quadrature and the azimuthal direction $\phi\in[0,\pi/2]$ is defined as symmetric with respect to $\phi=\pi/4$. The coordinates $\phi_{i}$ and weights $w_{i}$ are found by the solution of a nonlinear system of equations, which is solved by Newton iterations (Abu-Shumays, 1977) and the number of discrete ordinates per octant is $N_{\theta}\times N_{\phi}/2$, where $N_{\theta}$ is the order of the polar quadrature and $N_{\phi}$ order of the azimuthal quadrature. For a fixed order of quadrature, twice as many coordinates and weights are needed for the $\phi$ variable than for the $\theta$ variable, otherwise the quadrature would not be consistent (Abu-Shumays, 1977).
Using the above mentioned schemes we (Tres, 2015; Hu et al., 2016) investigated the property of exactly integrating highest order (direction cosines) polynomials. Such property is related to the error in the numerical approximation of the integral term, in Eq. (1). As a result, it was found that the square Legendre-Chebyshev quadrature ($P_{N}T_{N}$) showed a better performance on this aspect, in the sense of integrating exactly higher-order polynomials for a fixed quadrature order $N$. It is a very relevant issue when analyzing the convergence of the discrete ordinate solution according to Madsen’s theorem (Madsen, 1971), as we mentioned in this text. On the other hand, the $(QR)$ quadrature scheme has shown better performance (Barichello et al., 2016) to mitigate the so-called “ray effect” - a well-known numerical issue associated with the discrete-ordinates method in multidimensional geometries. Later, Spence (2015) proposed new approaches to generate ($QR$) schemes avoiding the ill-conditioning appearing in the previous formulation (Abu-Shumays, 1977). In Nenê (2020), a complete study and discussion on the generation of such schemes was developed, through the use of non-classical orthogonal polynomials.
The influence of the quadrature schemes on the solution of the transport/RTE equation has been investigated in several of our works (Da Cunha et al., 2015; Moura et al., 2018; Barichello et al., 2022). It is important to note that, while using the $LQ_{N}$ schemes we were able to consider only up to $36$ discrete directions per octant, in our simulations, the alternative schemes allowed us to consider, for example, up to $98$ (Barichello et al., 2022) and $253$ (Barichello et al., 2025). In fact, as the anisotropy of the medium increases, as the case of optical tomography applications, it is mandatory to increase the number of directions.
@@ 5.2 The ADO-Nodal Formulation
We write the two-dimensional discrete ordinates approximation of Eq. (1) in Cartesian geometry as (Barichello et al., 2022)
@@
for i=1,,Mi=1,...,M, where $M$ is the total number of discrete directions in the upper four octants of the unit sphere and $w_{n}$ are the associated weights, according to a chosen quadrature scheme; $\beta=\kappa+\sigma_{s}$ is the extinction coefficient, $\kappa$ is the absorption coefficient and $\sigma_{s}$ is the scattering coefficient; $S(x,y)$ is the source term; $p(\Omega_{n}\cdot\Omega_{i})$ is the phase function of scattering from the incoming direction $\Omega_{i}$ to the scattered direction $\Omega_{n}$. We note that at this point we are dealing with a system of first-order partial differential equations.
@@
We use a nodal method (Badruzaman, 1985) to solve Eq. (62), for each direction $i$, in the physical domain in the $(x,y)$ plane. To do this, the domain is divided into $r=1,\cdots,R$ regions (nodes), where each region $r$ is defined for $(x,y)\in[a_{h-1},a_{h}]\times[b_{k-1},b_{k}]$ (see Fig. (5)).
@@
Since we use the ADO formulation developed in Section 4 to solve the ODE system resulting from the application of the nodal scheme, we refer to this approach as the ADO-Nodal formulation (Barichello et al., 2017).
When deriving the two-dimensional model we assume that intensities are the same at $\bm{\Omega_{i}}=(\xi_{i},\mu_{i},\eta_{i})$ and $\bm{\Omega_{i}}=(-\xi_{i},\mu_{i},\eta_{i})$. It will be used to deal with the phase-function formulation (Barichello et al., 2025). In addition to that, when working with the angular coordinates on the plane, to obtain, for instance, the average angular intensity along the $y$ direction, we follow (Barichello et al., 2011) and start ordering the discrete directions (for $i=1,\ldots,M/2$) such that (Fig. (6))
@@
We refer to this ordering as $x$-scheme. Keeping all of that in mind,
Figure 6: Ordering Discrete Directions, $LQ_{4}$, $M=12$
@@
It is worth noting, regarding the integration performed on the variable $y$, that the terms $I_{r}\left(x,b_{k-1}^{r},\bm{\Omega}_{i}\right)$ and $I_{r}\left(x,b_{k}^{r},\bm{\Omega}_{i}\right)$ present in Eqs. (68), refer to the intensities on the boundaries of the regions. These terms may be known only in the incoming directions on the boundaries $y=0$ and $y=b$. In the exiting directions and at the boundaries of the internal nodes, these terms remain unknown. Thus, the resulting systems of equations for the average intensitites have more unknowns than equations, as usual in nodal schemes. It requires the use of auxiliary equations to close the system. In this work, the unknown intensities are approximated by constant functions. We have investigated this issue and proposed alternative approximations (Prolo Filho and Barichello, 2014; Cromianski et al., 2019) for the unknown (“leakage”) terms. It turns out the constant functions seem to be the best choice when balancing accuracy and computational cost.
We proceed similarly to derive the system for the average intensity along variable $x$. For more details see Barichello et al. (2022). It also requires an specific ordering for the discrete directions, we refer to as $y$-scheme, where, for $i=1,\ldots,M/2$,
@@
and after many algebraic manipulations and extensive derivation (Barichello et al., 2017, 2022, 2025) we obtain an eigenvalue problem
@@
The ADO method is a spectral method. The derivation of the eigenvalue problem is a key issue in this formulation. We note that with this derivation the eigenvalue problem derived for the ADO-Nodal formulation for two-dimensional problems is quite general. It is written in terms of the exact representation of the phase-function. Once one changes the transport theory field and the exact representation of the scattering pattern changes, the formulation is already developed. We note that in previous works (Barichello et al., 2022) the eigenvalue problem was derived based on the expansion of the phase-function for an arbitrary order of anisotropy $L$. In such case,
@@
$\zeta^{p}_{yl}(\Omega_{j})=\begin{cases}\textup{sen}(p\varphi_{j})\textup{sen}(p\varphi_{i})&\text{if p even }\\ \textup{cos}(p\varphi_{j})\textup{cos}(p\varphi_{i})&\text{if p odd }.\end{cases}$ (78)
@@
In this subsection, through the two-dimensional transport equation, we generated a system of one-dimensional equations, allowing the use of ADO method (Barichello and Siewert, 1999) for its solution. The homogeneous solution obtained by this method is constructed in terms of the solution of an eigensystem. It is relevant to stress that the ADO method when combined with the multidimensional quadratures studied preserves the property of reduction of the order of the associated eigenvalue problems thereby providing solutions with a higher degree of computational efficiency.
@@
$I_{yr}^{H}\left(x,\bm{\Omega}_{i}\right)=\sum_{j=1}^{M/2}\bigg{[}A_{j,r}\,\Phi_{yr}\left(\nu_{jr},\bm{\Omega}_{i}\right)e^{-\frac{(x-a_{h-1}^{r})}{\nu_{jr}}}+A_{j+\frac{M}{2},r}\,\Phi_{yr}\left(\nu_{jr},\bm{\Omega}_{i+\frac{M}{2}}\right)e^{-\frac{(a_{h}^{r}-x)}{\nu_{jr}}}\bigg{]}$ (81)
$I_{yr}^{H}\left(x,\bm{\Omega}_{i+\frac{M}{2}}\right)=\sum_{j=1}^{M/2}\bigg{[}A_{j,r}\,\Phi_{jr}\left(\nu_{jr},\bm{\Omega}_{i+\frac{M}{2}}\right)e^{-\frac{(x-a_{h-1}^{r})}{\nu_{jr}}}+A_{j+\frac{M}{2},r}\,\Phi_{yr}\left(\nu_{jr},\bm{\Omega}_{i}\right)e^{-\frac{(a_{h}^{r}-x)}{\nu_{jr}}}\bigg{]}$ (82)
@@
The general solution to the problem is obtained by adding the homogeneous solution to the particular solution. To propose the particular solution, it is first necessary to define the source terms as in Eqs.(65), (66), which include unknown intensities. Defining these components requires the use of auxiliary equations. According to previous works (Barichello et al., 2017; Cromianski et al., 2019; Barichello et al., 2022), a good choice is defining these auxiliary equations by assuming the unknown terms as constant functions, as we mention earlier in this text.
@@
Defining the unknown intensities at the boundaries of the regions as constants leads to the choice of constant approximations for the particular solution, that is,
@@
Thus, we can write the general solution for the average intensity along the $y$ direction, by adding the homogeneous (Eqs.(81) and (82)) and particular solution (Eq. (85)), taking into account the defined ordering of directions ($y$-scheme), for $i=1,\ldots,M/2$, as
@@
For developing the formulation for the average intensity in the $x$ direction, we follow similar procedure Barichello et al. (2022) and express the general solution , assuming the previously defined ordering directions ($x$-scheme), in the form
@@
To completely establish the average intensities in the variables $x$ and $y$ we need to solve a $4M(HK)$ linear system for the arbitrary coefficients ($A_{j}$, $B_{j}$, $Z_{j}$ and $W_{j}$, with $j=1,\ldots,M$). The order of the linear system is defined by the number of discrete directions as well as the number of the spatial nodes. High-quality solutions are expected as both, the number of discrete directions and the refinement of the spatial mesh, increase. We have investigated he performance of direct and iterative methods, for the solution of the linear systems, along with domain decomposition techniques and parallel implementation (Da Cunha et al., 2015; Rui et al., 2021; Barichello and da Cunha, 2023). Alternative arrangements in the configuration of the equations allowed solutions to higher order systems. A dependence on the type of the quadrature scheme as well as the class of problems to be solved (neutron or radiation problems, for instance) directly affect the final choice of the numerical algorithm (Barichello and da Cunha, 2023; Moura et al., 2018). Iterative methods such as GMRES, Loose GMRES and TFQMR were successful in neutron applications, for different choices of quadrature schemes. For the radiative transfer problem discussed here, those methods were unable to solve the linear system and the alternative was to try to obtain its solution via direct methods (LU factorization). Due to its high sparsity, which lead to unacceptable levels of fill-in the factorization, a reordering of the equations of the system was necessary to allow an application of a sparse block diagonal subroutine (Rui et al., 2021; Barichello et al., 2022; Barichello and da Cunha, 2023).
6 The Integrodifferential Formulation: Neutrons
An interesting feature of BE is its applicability in different areas. Specific literature is usually found for each of these applications. Due to inherent features of the radiation interaction with the media there are differences on the nomenclature and quantities of interest for each case. Although the ADO-Nodal formulation was developed in the previous section we find interesting to stress particular aspects of the model and some issues we investigated in the context of the two-dimensional neutron transport equation.
We begin with the time-independent neutron transport equation which considers the distribution of the particles in non-multiplying homogeneous media, with one energy group, written as follows

### v2

@@ 4.1 Numerical Results and Contributions
The ADO formulation for problems in one-dimensional geometry was developed in this section in the context of the RTE. In addition to the works already cited, the methodology has been applied to a wide range of problems in various areas of transport theory by different authors. It is worth listing here the treatment of models which include polarization effects [16] of interest in astrophysics, and that, according to our knowledge, has not been solved before; the challenging searchlight problem [18]; ADO for different geometries [13]; medical applications [39]; the key issues of criticality [34] and safeguards [11] in nuclear engineering; to mention just a few. Later in this work, applications in the field of rarefied gas dynamics will be discussed. We do not list numerical results generated by the solution here. They can be checked in the listed references. However, a central point to note is that in all these applications, the ADO solution proved to be concise and accurate, requiring a lower computational cost compared to what was available in the literature. The natural challenge, considering more realistic problems, was to extend the idea to multidimensional problems, which are the subject of the next section.
@@ 5 The ADO Solution for Two-dimensional Geometry Problems
Purely numerical methods dominate simulations for multidimensional transport models. The use of analytical techniques is limited and very challenging. However, they are of great importance for generating benchmark solutions. One of the first issues regarding discrete ordinate methods, which already differ greatly from one-dimensional problems, is the definition of numerical quadrature schemes for integration on the unit sphere. Madsen [47] identified the truncation error in the approximation of the integral term of the equation as a fundamental point in the convergence of the discrete-ordinate solution.
@@ 5.1 Numerical Quadrature Schemes
The choice of an appropriate quadrature scheme to approximate the integral term in Eq. (1) is a relevant issue, in particular for multidimensional problems. Traditionally, the quadratures were required to preserve the reflective and rotational symmetry properties and, in this context, the classical Level Symmetric ($LQ_{N}$) quadrature scheme [43] is widely used for dealing with multidimensional discrete ordinates problems (see Fig. 3). It is well known, however, that such a scheme is limited to order $N\leq 20$. The number of directions per octant is given by
$M=\frac{N(N+2)}{8}.$
In our studies, we have additionally considered triangular ($P_{N}T_{N}S_{N}$) and quadrangular ($P_{N}T_{N}$) Legendre-Chebyshev schemes [44, 26] and Quadruple Range ($QR$) schemes [1], which enable the generation of numerical results for much higher quadrature orders, even if some (rotational) symmetry properties are relaxed.
The Legendre-Chebyshev quadrature schemes use a one-dimensional Gaussian quadrature to treat the polar variable $\theta$ and Chebyshev quadrature for the azimuthal variable $\phi$ [44, 26]. Thus, in a three-dimensional geometry, the ordinates $\xi=\cos\theta$ are given by the roots of the Legendre polynomials ($P_{N}$). The azimuthal angular variable in each level is defined by the roots of the Chebyshev polynomials ($T_{N}$) of the first kind. The Legendre-Chebyshev quadrangular scheme ($P_{N}T_{N}$) [26] defines $M=(N^{2}/4)$ discrete directions per octant, where for each level $\xi_{i}$ the corresponding directions have the same weight (Fig. 4). For the two-dimensional case, one coordinate is disregarded. Still, the corresponding quadrature weights are given in terms of $w_{i}$, the weights of the Gauss-Legendre quadrature.
Figure 4: Legendre-Chebyshev Quadrangular Scheme
For the Legendre-Chebyshev Triangular scheme ($P_{N}T_{N}S_{N}$) [44], on the other hand, the $M=N(N+2)/8$ discrete directions per octant follow a pattern analogous to the Level Symmetric quadrature scheme. The set of azimuthal angles for each level $\xi_{i}$ is defined as
@@
The Quadruple Range quadrature ($QR$) scheme, proposed by Abu-Shumays [1], is a product quadrature. As a general description, the polar angle $\theta$ is approximated by Gauss-Christoffel quadrature and the azimuthal direction $\phi\in[0,\pi/2]$ is defined as symmetric with respect to $\phi=\pi/4$. The coordinates $\phi_{i}$ and weights $w_{i}$ are found by the solution of a nonlinear system of equations, which is solved by Newton iterations [1] and the number of discrete ordinates per octant is $N_{\theta}\times N_{\phi}/2$, where $N_{\theta}$ is the order of the polar quadrature and $N_{\phi}$ order of the azimuthal quadrature. For a fixed order of quadrature, twice as many coordinates and weights are needed for the $\phi$ variable than for the $\theta$ variable, otherwise the quadrature would not be consistent [1].
Using the above mentioned schemes we [21, 64, 41] investigated the property of exactly integrating the highest order (direction cosines) polynomials. Such property is related to the error in the numerical approximation of the integral term, in Eq. (1). As a result, it was found that the square Legendre-Chebyshev quadrature ($P_{N}T_{N}$) showed a better performance on this aspect, in the sense of integrating exactly higher-order polynomials for a fixed quadrature order $N$. It is a very relevant issue when analyzing the convergence of the discrete ordinate solution according to Madsen’s theorem [47], as we mentioned earlier in this text. On the other hand, the $(QR)$ quadrature scheme has shown better performance [21] to mitigate the so-called “ray effect” - a well-known numerical issue associated with the discrete-ordinates method in multidimensional geometries. Later, Spence [62] proposed new approaches to generate ($QR$) schemes avoiding the ill-conditioning appearing in the previous formulation [1]. In [49], a complete study and discussion on the generation of the (QR) schemes was developed, through the use of non-classical orthogonal polynomials.
The influence of the quadrature schemes on the solution of the transport/RTE equation has been investigated in several of our works [35, 48, 15]. It is important to note that, while using the $LQ_{N}$ schemes, we were able to consider only up to $36$ discrete directions per octant in our simulations. The alternative schemes allowed us to consider, for example, up to $98$ [15], and $253$ [9]. In fact, as the anisotropy of the medium increases, as in the case of optical tomography applications, we found it mandatory to increase the number of directions.
@@ 5.2 The ADO-Nodal Formulation
We write the two-dimensional discrete ordinates approximation of Eq. (1) in Cartesian geometry as [9, 15]
@@
for i=1,,Mi=1,...,M, where $M$ is the total number of discrete directions in the upper four octants of the unit sphere and $w_{n}$ are the associated weights, according to a chosen quadrature scheme; $\beta=\kappa+\sigma_{s}$ is the extinction coefficient, $\kappa$ is the absorption coefficient and $\sigma_{s}$ is the scattering coefficient; $S(x,y)$ is the source term; $p(\Omega_{n}\cdot\Omega_{i})$ is the phase function of scattering from the incoming direction $\Omega_{i}$ to the scattered direction $\Omega_{n}$. We note that at this point, we are dealing with a system of first-order partial differential equations.
We use a nodal method [3] to solve Eq. (62), for each direction $i$, in the physical domain in the $(x,y)$ plane. To do this, the domain is divided into $r=1,\cdots,R$ regions (nodes), where each region $r$ is defined for $(x,y)\in[a_{h-1},a_{h}]\times[b_{k-1},b_{k}]$ (see Fig. 5).
@@
Since we use the ADO formulation developed in Section 4 to solve the ODE system resulting from the application of the nodal scheme, we refer to this approach as the ADO-Nodal formulation [12].
When deriving the two-dimensional model, we assume that intensities are the same at $\bm{\Omega_{i}}=(\xi_{i},\mu_{i},\eta_{i})$ and $\bm{\Omega_{i}}=(-\xi_{i},\mu_{i},\eta_{i})$. It will be used to deal with the phase-function formulation [9]. In addition to that, when working with the angular coordinates on the plane, to obtain, for instance, the average angular intensity along the $y$ direction, we follow [6] and start ordering the discrete directions (for $i=1,\ldots,M/2$) such that (Fig. 6)
Figure 6: Ordering Discrete Directions, $LQ_{4}$, $M=12$
@@
It is worth noting, regarding the integration performed on the variable $y$, that the terms $I_{r}\left(x,b_{k-1}^{r},\bm{\Omega}_{i}\right)$ and $I_{r}\left(x,b_{k}^{r},\bm{\Omega}_{i}\right)$ present in Eqs. (68), refer to the intensities on the boundaries of the regions. These terms may be known only in the incoming directions on the boundaries $y=0$ and $y=b$. In the existing directions and at the boundaries of the internal nodes, these terms remain unknown. Thus, the resulting systems of equations for the average intensities have more unknowns than equations, as usual in nodal schemes. It requires the use of auxiliary equations to close the system. In this work, the unknown intensities are approximated by constant functions. We have investigated this issue and proposed alternative approximations [54, 33] for the unknown (“leakage”) terms. It turns out the constant functions seem to be the best choice when balancing accuracy and computational cost.
We proceed similarly to derive the system for the average intensity along variable $x$. For more details, see [15]. It also requires a specific ordering for the discrete directions, we refer to as $y$-scheme, where, for $i=1,\ldots,M/2$,
@@
and after many algebraic manipulations and extensive derivations [12, 15, 9] we obtain an eigenvalue problem
@@
The ADO method is a spectral method. The derivation of the eigenvalue problem is a key issue in this formulation. We note that with this derivation, the eigenvalue problem derived for the ADO-Nodal formulation for two-dimensional problems is quite general. It is written in terms of the exact representation of the phase-function. Once one changes the transport theory field and the exact representation of the scattering pattern changes, the formulation is already developed. We note that in previous works [57, 15] the eigenvalue problem was derived based on the expansion of the phase-function for an arbitrary order of anisotropy $L$. In such case,
@@
$\zeta^{p}_{yl}(\Omega_{j})=\begin{cases}\textup{sen}(p\varphi_{j})\textup{sen}(p\varphi_{i})&\text{if p even }\\ \textup{cos}(p\varphi_{j})\textup{cos}(p\varphi_{i})&\text{if p odd }\end{cases}$ (78)
@@
In this subsection, through the two-dimensional transport equation, we generated a system of one-dimensional equations, allowing the use of ADO method [17] for its solution. The homogeneous solution obtained by this method is constructed in terms of the solution of an eigensystem. It is relevant to stress that the ADO method, when combined with the multidimensional quadratures studied, preserves the property of reduction of the order of the associated eigenvalue problems, thereby providing solutions with a higher degree of computational efficiency.
@@
$I_{yr}^{H}\left(x,\bm{\Omega}_{i}\right)=\sum_{j=1}^{M/2}\bigg[A_{j,r}\,\Phi_{yr}\left(\nu_{jr},\bm{\Omega}_{i}\right)e^{-\frac{(x-a_{h-1}^{r})}{\nu_{jr}}}+A_{j+\frac{M}{2},r}\,\Phi_{yr}\left(\nu_{jr},\bm{\Omega}_{i+\frac{M}{2}}\right)e^{-\frac{(a_{h}^{r}-x)}{\nu_{jr}}}\bigg],$ (81)
$I_{yr}^{H}\left(x,\bm{\Omega}_{i+\frac{M}{2}}\right)=\sum_{j=1}^{M/2}\bigg[A_{j,r}\,\Phi_{jr}\left(\nu_{jr},\bm{\Omega}_{i+\frac{M}{2}}\right)e^{-\frac{(x-a_{h-1}^{r})}{\nu_{jr}}}+A_{j+\frac{M}{2},r}\,\Phi_{yr}\left(\nu_{jr},\bm{\Omega}_{i}\right)e^{-\frac{(a_{h}^{r}-x)}{\nu_{jr}}}\bigg],$ (82)
@@
The general solution to the problem is obtained by adding the homogeneous solution to the particular solution. To propose the particular solution, it is first necessary to define the source terms as in Eqs. (65), (66), which include unknown intensities. Defining these components requires the use of auxiliary equations. According to previous works [12, 33, 15], a good choice is defining these auxiliary equations by assuming the unknown terms as constant functions, as we mention earlier in this text.
@@
Defining the unknown intensities at the boundaries of the regions as constants as well as $S_{yr}(x)$ in Eq. (68), allows us the choice of constant approximations for the particular solution, that is,
@@
Thus, we can write the general solution for the average intensity along the $y$ direction, by adding the homogeneous (Eqs. (81) and (82)) and particular solution (Eq. (85)), taking into account the defined ordering of directions ($x$-scheme), for $i=1,\ldots,M/2$, as
@@
For developing the formulation for the average intensity in the $x$ direction, we follow a similar procedure [15] and express the general solution, assuming the previously defined ordering directions ($y$-scheme), in the form
@@
To completely establish the average intensities in the variables $x$ and $y$ we need to solve a $4M(HK)$ linear system that results to be only for the arbitrary coefficients ($A_{j}$, $B_{j}$, $Z_{j}$, and $W_{j}$, with $j=1,\ldots,M$). The order of the linear system is defined by the number of discrete directions as well as the number of the spatial nodes. High-quality solutions are expected as both, the number of discrete directions and the refinement of the spatial mesh increase. We have investigated the performance of direct and iterative methods, for the solution of such linear systems, along with domain decomposition techniques and parallel implementation [35, 56, 8]. Alternative arrangements in the configuration of the equations allowed solutions to higher-order systems. A dependence on the type of the quadrature scheme as well as the class of problems to be solved (neutron or radiation problems, for instance) directly affects the final choice of the numerical algorithm [8, 48]. Iterative methods such as GMRES, Loose GMRES, and TFQMR were successful in neutron applications for different choices of quadrature schemes. For the thermal radiation problems [15] those methods were unable to satisfactorily solve the linear system. The alternative was to obtain its solution via direct methods (LU factorization). Due to its high sparsity, which leads to unacceptable levels of fill-in the factorization, a reordering of the equations of the system was necessary to allow an application of a sparse block diagonal subroutine [56, 15, 8].
6 Neutron Transport
An interesting aspect of BE is its applicability in different areas. Specific literature is usually found for each of these applications. Due to inherent features of the radiation interaction with the media, there are differences on the nomenclature and quantities of interest for each case. While the ADO-Nodal formulation was developed in the previous sections in the context of radiative transfer applications, we find interesting to stress particular aspects of the model and some issues we investigated in the context of the two-dimensional neutron transport equation.
We begin with the time-independent neutron transport equation, which considers the distribution of the particles in non-multiplying homogeneous media, with one energy group, written as follows